OpenStudy (angiepangie0726):

What is the equation, in vertex form, of the quadratic function that has a vertex at V(−5, −4) and passes through point P(1, 8)?

OpenStudy (welshfella):

the general vertex form is a(x - b^2 + c where the vertex is at (b,c) and the value of a can be found if you know one point on the curve

OpenStudy (welshfella):

a(x - b^2 + c = 0

OpenStudy (angiepangie0726):

a(x-(-5)^2+(-4)=0

OpenStudy (welshfella):

so here b = -5 and c = -4 so we can write a(x - (-5)) + (-4) = 0

OpenStudy (welshfella):

right which simoplifies to a(x + 5)^2 - 4 = 0 now plug in the point (1,8) and solve for a and you have the answer

OpenStudy (welshfella):

I should have written it as y = a(x + 5)^2 - 4 so plug in y = 8 and x = 1 ( I'm not at my best today!)

OpenStudy (welshfella):

8 = a(1+ 5)^2 - 4 solve for a

OpenStudy (angiepangie0726):

Ok let me write this down and I will let you know what I get :)

OpenStudy (angiepangie0726):

4

OpenStudy (angiepangie0726):

y equals start fraction one over three end fraction left parenthesis x plus five right parenthesis squared minus four y = 3(x + 5)2 − 8 y equals start fraction negative one over three end fraction left parenthesis x plus five right parenthesis squared plus four y equals start fraction one over three end fraction left parenthesis x minus five right parenthesis squared minus four

OpenStudy (angiepangie0726):

These are my choices though:/

OpenStudy (welshfella):

No 6^2 *a = 12 a = 12/36 = 1/3

OpenStudy (angiepangie0726):

ohh

OpenStudy (angiepangie0726):

So it cant be B or C

OpenStudy (welshfella):

so its y = (1/3)(x + 5)^2 - 4

OpenStudy (welshfella):

right

OpenStudy (welshfella):

which one is it A or D?

OpenStudy (angiepangie0726):

Thank you so much! I have one more question, Do you think you can help me out?? And it would be A

OpenStudy (welshfella):

A is correct Please re post another question.